# Conceptualisation, Methodology, Software, Validation, Writing - original draft

### v2

@@ Abstract
To correct the outdated and incomplete topology of low-voltage distribution networks (LVDNs) solely based on voltage magnitudes, a data-driven approach is developed based on machine learning algorithms and correlation analysis. Meanwhile, to address the similarity among smart meter (SM) data induced by distributed photovoltaic (PV) systems, a time-based SM data selection strategy is combined with the proposed correlation analysis. Unlike offline approaches, the proposed approach uses up-to-date voltage magnitudes to help distribution system operators determine switch states via supervised learning and refine user–feeder connections and customer phase labels through a modified hierarchical clustering algorithm. The feasibility and robustness of the proposed approach are validated using modified real-world LVDNs and multiple incomplete SM datasets collected from customers in the Netherlands. The results demonstrate that the time-based SM data selection strategy effectively mitigates its impact on phase identification, and the corrected topology not only improves network observability but also supports network operators in load balancing and PV consumption.
@@ Introduction
Balanced distribution network contributes to reducing voltage violations, line congestion, power loss, etc. Most existing phase balancing and reconfiguration problems are formulated as mixed-integer optimization problems that depend on network topologies [10098964]. However, these topologies are often unavailable or inaccurate for distribution system operators (DSOs) due to outdated or missing data, such as limited phase labels [vanin2024phase]. The increasing uncertainty of distributed energy resources (DERs), including household photovoltaic (PV) systems, heating pumps etc., impacts the topology reconfiguration frequency and challenges the identification of low-voltage distribution network (LVDN) topology [10026490, 10347462, 10475702]. Moreover, existing topology identification methods rely on complete smart meter (SM) datasets, the availability of which is often limited due to privacy concerns [9696306, costa2022identification]. Synthetic networks and benchmark models presented in [birchfield2016grid, 2020Non] are insufficient to represent the diversity of European LVDNs for practical use by DSOs (e.g., state estimation). Thus, practical topology correction methods for LVDNs are needed for active management and DERs penetration.
The topology identification problem in distribution networks (DNs) could be briefly classified into three categories based on the main issues they are aiming to solve: (1) switch state identification, (2) connection lines and their length identification, and (3) phase identification. The switch state identification was modeled as a classification problem in [9758816], solved by a trained deep neural network and providing quick estimations under the selected features. To recover accurate connection lines between transformers and customers, a probabilistic graphical model-based method was developed in [8456535] by restricting the propagation of SM data error. A power band-based strategy is proposed for dataset selection, aiming to enhance the accuracy of phase identification and transformer-to-meter matching [10533846]. In [liu2023hybrid], a hybrid data-driven approach integrating a partial correlation analysis strategy and a linear regression model is proposed to estimate topology from limited SM data. A similar work in [wang2023data] established a data-driven model based on the impedance matrix to correct topology and identify the impedance of connection lines from SM data.
Based on an alternating direction method of multipliers, a robust topology identification method is proposed to estimate the line connection and their parameters using $\mu$-PMUs and SM datasets [9729419]. Based on incomplete and asynchronous SM data, an optimization-based data-driven approach is constructed to recover the approximate topologies in LVDNs with open datasets [liu2025topology]. A two-step approach is introduced to identify the topology of LVDNs with multiple meter types, which provide different resolution datasets, mitigating the dependence on the SMs [10400804]. For networks with low observability, a distance matrix is introduced in [10612997] to correct the topology and is also effective in detecting latent nodes positioned between observed nodes along network paths (i.e., intermediate latent nodes rather than endpoints). A physics-informed neural network (PINN) is constructed to estimate the topology by representing the nodes without meters, whose parameters are adjusted by a reinforcement learning-based approach [10138375]. Although the above methods generate accurate topologies of DNs, the requirements of complete SM data and prior topology information (e.g., the topology candidates) make them infeasible in practice due to data privacy-related concerns. Moreover, optimizing the line resistance and reactance as independent variables may lead to unrealistic estimation, as the properties are inherently interdependent and influenced by physical environments, such as temperature, humidity.
Related to phase identification, current approaches rely on voltage magnitude data and demand profiles [7098424, garcia2025data]. Since the correlation of time-series voltage magnitude data from the same phase is stronger than that from different phases, correlation analysis is normally used. Given voltage magnitude datasets, spectral clustering and $k$-means clustering approaches are adopted to identify customer phase label [8955915]. The mixed integer linear programming model in [9409966] is constructed to identify the phase labels of single-phase customers by minimizing the gap between transformers’ measurements and the estimated data (e.g., power and voltage magnitude) in each phase. An optimization model-based approach was introduced in [10098964] to estimate customer phase labels in LVDNs with limited available recordings. Saliency analysis (SA) was integrated to extract the underlying features aiming to enhance the accuracy of phase label identification [7604110] SA and statistical tests are combined in [jimenez2020phase] to guarantee the identification accuracy of the data pairs with weak correlations. Voltage magnitude fluctuations among customers connected to different phases may exhibit similarities due to variations induced by the medium-voltage (MV) network or similar load patterns. To mitigate this effect, a high-pass filter is applied during time-series data pre-processing in the phase identification approach [hosseini2020machine]. Beyond hard classification (i.e., assigning customers a binary label of 0 or 1), soft classification is also considered for phase identification. The method in [garcia2023phase] assigns three probability values to customers. The highest probability indicates the most likely phase label for the user.
@@
MV/LV Medium/Low voltage LVDN Low-voltage distribution network DNs Distribution networks SM Smart meter DSO Distribution system operator DERs Distribution energy resources SA Saliency analysis PV Photovoltaic EV Electical vehicles HC Hierarchical clustering algorithm RF Random forest algorithm KNN K-nearest neighbors algorithm GNB Gaussian Naive Bayes algorithm SVM Support vector machine algorithm ANN Artificial neural network PINN Physics-informed neural network
@@
        PCC Pearson correlation coefficient         MFP Modified PCC based Fisher z-transformation         GIS Geographic Information System         CN1 Network with two complex-structured feeders         CN2 Network with one simple- and one complex-structured feeder         SN Network with two simple-structured feeders         SNL Network with two longer simple-structured feeders         SNB Network with two simple-structured feeders with more branches         CC_PV CN1 with PV at each house         SC_PV SN with PV at each house         $C$ Complex feeder from CN1         SC Simple feeder from SN         CC_3_NU Complex feeder with three-phase non-uniform loads         SC_3_NU Simple feeder with three-phase non-uniform loads
However, the key step in the above approaches is the stepwise subtraction of the identified customers or users with the highest possibility labels, which ensures the accuracy of the phase labels of the customers who are located far away from the substation. Nevertheless, this key step might induce inaccuracies, since the accuracy of the current step is subject to that of the previous one. To address this issue, based on a modified genetic algorithm, a novel data-driven phase identification approach is introduced to identify the phase labels under incomplete datasets [jimenez2021new]. Based on multiple linear regression models, [cunha2020automated] presents a hybrid topology identification method to simultaneously identify line parameters and phase connection from raw SM data. Nevertheless, the underlying assumptions of the aforementioned methodologies—such as the presumption of complete datasets devoid of SM errors, and the omission of MV network impacts—significantly constrain their practical applicability [therrien2021assessment]. Fig. 1 presents several underground feeder deployment patterns in the Netherlands [web1], which indicates the challenges of LVDN topology identification and correction in practice, e.g., multiple cables under the same street. In addition, some of the above approaches are sensitive to missing data and measurement errors, e.g., SA-based approaches. Furthermore, the user-feeder connection identification issue is rarely considered in the aforementioned papers due to the assumption that one main feeder connects to the transformer.
@@
To overcome these limitations, we propose a data-driven approach for topology correction in LVDNs using incomplete voltage data. Table 1 highlights the key differences between the proposed approach and existing approaches. By relying solely on voltage magnitude data, this approach preserves privacy and minimizes SM updates, as it releases the need for phase angle measurements from households. The corrected topology provides the foundational network visibility needed to implement load balancing, facilitate DERs integration, and alleviate LV line congestion, enabling active management. The main contributions are summarized as follows:
@@
  • A data-driven topology correction approach is proposed to correct the missed and outdated recordings in LVDN topologies by three steps: (1) switch state identification, (2) user-feeder correction, and (3) phase label identification. The proposed approach relies solely on voltage magnitude data, reducing privacy risks. Compared to previous methods that depend on complete SM datasets, this approach allows the DSOs to correct the topologies using limited SM data with noise from incomplete household measurements (i.e., not all house SM is accessible).
@@
  • The operational states of the switch bars (including multiple switchgear) are encoded using Label Encoding rules, transforming the independent identification task of each switchgear into a classification problem that can be effectively addressed using existing supervised learning methods. In correlation analysis, a modified distance function is integrated into traditional Hierarchical Clustering (HC) to amplify the distinction between loads that are geographically close but belong to different phases or cables, thereby improving identification accuracy.
  • To mitigate the impact of PV power on voltage-magnitude based correlation analysis, we propose a time-based SM data selection strategy to extract voltage magnitude measurements primarily influenced by household loads (e.g., during nighttime or low-PV-output periods). The resulting datasets enhance topology correction accuracy, particularly in scenarios with noisy or incomplete datasets.
@@
Table 1: Input data and outputs in topology problems
Method Ref. Input Data Robustness Output
$V$ $P/I$ Limited GIS DERs SM Missed Switch user-feeder Phase
topology data impact errors data state connections labels
DNN [9758816] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
[10138375] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$
Linear regression [8456535] $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
[cunha2020automated] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
Optimization based [liu2025topology] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$
[9409966] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$
clustering-based [8955915] $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$
[10533846] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$
Heuristic Algorithms [jimenez2021new] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$
Our Method $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@
The proposed topology correction approach is illustrated in Fig. 2, which is composed of three steps: (1) switch state identification, (2) user-feeder connection identification and (3) phase label identification. The input includes customer voltage magnitude and incomplete recordings of the user-feeder connection, and the output is the corrected topology with accurate customer’s user feeder connections and phase labels, including single-phase and three-phase customers.
@@
Figure 3: Illustrative example of switch deployment and states in LVDNs: 1 represents the activated switch and vice versa [abb2025].
@@
    $\hat{y}=\text{argmax}_{j}\sum_{b=1}^{B}\mathbb{I}(h_{b}(x)=j),$ (5)
where $h_{b}(x)$ is the prediction of the $b$-th tree and $\mathbb{I}$ is the indicator function. Given a voltage sample $V_{i}$, each trained decision tree independently predicts $h_{b}(x)$ and the final labels of the switch are obtained by the majority voting mechanism.
@@
Given the switch state obtained from Section 2.1, the feeder-transformer connection is available, and the number of feeders connected to the same transformer is assumed to be accessible for the DSO and taken as the number of clusters (i.e., $K_{s}$). Lines 1-13 in Algorithm 1 are to obtain the MFP of the input voltage magnitude dataset. Lines 16-23 are to find the link matrix of matrix $D$. From line 23 onwards, input data are clustered into a $K_{s}$ cluster using the obtained linkage matrix. If two users share the same label, they are connected to the same underground feeder, and vice versa.
@@
$\hat{y}=\arg\max_{c}\sum_{n\in\mathcal{N}_{k}}\mathbb{I}(y_{n}=c)$ (10)
@@
Although various signal processing techniques, such as SA and high-pass filters, can be employed to mitigate high-frequency or low-frequency noise induced by household PV, the determination of appropriate parameters—such as cutoff frequency—must be tailored to specific datasets. This requirement presents practical challenges in implementing these methods effectively. Inspired by the power-band based strategies in [10533846], a time-based data selection strategy is introduced to enhance identification accuracy by screening summer data, especially for phase identification. The proposed time-based data selection strategy is easily deployed and fast. Two strategies for selecting relevant data are introduced:
@@
By applying these time based selection strategies, the impact of PV-induced voltage fluctuations can be mitigated, thereby improving the robustness of correlation-based phase identification methods.
@@
In this section, the feasibility and robustness of the proposed topology correction framework are validated using modified real-world LVDNs and SM data collected from customers in the Netherlands. The code used in this paper is available online  [^1] https://github.com/distributionnetworksTUDelft/Data-driven-approach-for-topology-correction-in-LVDNs-. The resolution of the dataset is 15 minutes, resulting in 96 data points per day. The base three-phase voltage level is 0.4 kV, and the loads of three-phase customers are non-uniformly distributed across Phases A, B, and C. The time thresholds $t_{1}$ and $t_{2}$ are set to 20 and 88. The voltage magnitude profiles are generated using a power flow model, being solved by using the Power Grid Model package [PowerGridModel].
@@
Figure 7: Switch state identification accuracy under incomplete dataset by: a) RF and b) SVM
Table 2: Accuracy (%) comparison under multiple scenarios
@@
Given datasets with random missed data points, the accuracy of RF and SVM were assessed, and the average accuracy under five simulations is shown in Fig. 7. In general, the accuracy of the two approaches improves significantly as the number of available SMs increases and slightly decreases as the missing percentage increases, demonstrating the robustness of the proposed method with sufficient data availability. When only 2 SMs data are available, the accuracy remains consistently low (around 55–60%), indicating the method’s limited capability under sparse data conditions. Besides, the sensitivity of SVM accuracy to the available data is less than that of RF, since the accuracy of SVM remains close when the number of available SM is larger than 10. These findings highlight the robustness of SVM and RF for switch state identification tasks in LVDNs.
@@
Table 3: Accuracy (%) comparison under multiple scenarios

### v3

@@ Abstract
To correct the outdated and incomplete topology of low-voltage distribution networks (LVDNs) using solely voltage magnitudes, a data-driven approach is developed by integrating machine learning algorithms with correlation analysis. Unlike conventional offline approaches, the proposed method leverages up-to-date voltage magnitude measurements within a unified framework, enabling the same dataset to be consistently utilized across all processing stages without repeated data preprocessing. Specifically, switch states are identified via supervised learning, while user–feeder connections and customer phase labels are refined using a modified hierarchical clustering algorithm. To address the similarity among smart meter (SM) data induced by distributed photovoltaic (PV) systems, a time-based SM data selection strategy is incorporated into the correlation analysis. The feasibility and robustness of the proposed approach are validated using modified real-world LVDNs and multiple incomplete SM datasets collected from customers in the Netherlands. The results demonstrate that the time-based SM data selection strategy effectively mitigates the impact of PV-induced similarity on phase identification, and that the corrected topology not only improves network observability but also supports distribution system operators in load balancing and PV integration.
@@ Introduction
Most existing phase balancing and topology reconfiguration problems are formulated as mixed-integer optimization problems that depend on network topologies [10098964, 11017695, 10571996]. However, these topologies are often inaccurate and outdated for distribution system operators (DSOs) due to missing recordings, topology maintenance and reconfiguration, such as congestion management  [vanin2024phase]. Thus, the topology of the low-voltage distribution network (LVDN) needs to be checked and corrected when it is outdated. The increasing uncertainty of distributed energy resources (DERs), including household photovoltaic (PV), heating pumps, etc., impacts the frequency of topology reconfiguration and challenges the correction of the low-voltage distribution network topology [10026490, 10347462, 10475702]. Moreover, the available smart meter (SM) datasets are often limited due to privacy concerns and random communication channel failure, challenging the topology correction [9696306, costa2022identification, dande2025consumer]. Synthetic European networks and benchmark models presented in [birchfield2016grid, 2020Non] are benchmarks for research but insufficient to represent the diversity of European LVDNs for practical use by DSOs (e.g., state estimation). Thus, practical topology identification and correction approaches are required for real-time topology updating for active management of LVDNs.
Literature review
The topology identification problem in distribution networks (DNs) could be briefly classified into three categories: (1) switch state identification, (2) connection lines and their parameters identification, and (3) phase identification. The switch state identification was taken as a classification problem in [9758816], solved by a trained deep neural network and providing quick estimations under the selected features. To recover accurate connection lines between transformers and customers, a probabilistic graphical model-based method was developed in [8456535] by restricting the propagation of SM data error. In [liu2023hybrid], a hybrid data-driven approach integrating a partial correlation analysis strategy and a linear regression model is proposed to estimate topology from limited SM data. Based on an alternating direction method of multipliers, a robust topology identification method is proposed to estimate the line connection and their parameters using $\mu$-PMUs and SM datasets [9729419]. A two-step approach is introduced to identify the topology of LVDNs with multiple meter types, which provide different resolution datasets, mitigating the dependence on the SMs [10400804]. Related to phase identification, current approaches rely on voltage magnitude data and demand profiles [garcia2025data]. Since the correlation of voltage magnitude data from the same phase is stronger than that from different phases, correlation analysis is normally used. Given voltage magnitude datasets, spectral clustering and $k$-means clustering approaches are adopted to identify customer phase label [8955915]. The mixed integer linear programming model in [9409966] is constructed to identify the phase labels of single-phase customers by minimizing the gap between transformers’ measurements and the estimated data (e.g., power and voltage magnitude) in each phase. Saliency analysis (SA) and statistical tests are combined in [jimenez2020phase] to guarantee the identification accuracy of the data pairs with weak correlations. Fig. 1 presents several underground feeder deployment patterns in the Netherlands [web1], which indicates the challenges of LVDN topology identification and correction in practice, e.g., multiple cables under the same street. In addition, some of the above approaches are sensitive to missing data and measurement errors, e.g., SA-based approaches. Furthermore, the user-feeder connection identification issue is rarely considered in the aforementioned papers due to the assumption that one main feeder connects to the transformer.
@@
MV/LV Medium/Low voltage LVDN Low-voltage distribution network DNs Distribution networks SM Smart meter DSO Distribution system operator DERs Distribution energy resources SA Saliency analysis PV Photovoltaic EV Electical vehicles HC Hierarchical clustering algorithm RF Random forest algorithm KNN K-nearest neighbors algorithm GNB Gaussian Naive Bayes algorithm SVM Support vector machine algorithm ANN Artificial neural network PCC Pearson correlation coefficient
@@
        MFP Modified PCC based Fisher z-transformation         GIS Geographic Information System         CN1 Network with two complex-structured feeders         CN2 Network with one simple- and one complex-structured feeder         SN Network with two simple-structured feeders         SNL Network with two longer simple-structured feeders         SNB Network with two simple-structured feeders with more branches         CC_PV CN1 with PV at each house         SC_PV SN with PV at each house         $C$ Complex feeder from CN1         SC Simple feeder from SN         CC_3_NU Complex feeder with three-phase non-uniform loads         SC_3_NU Simple feeder with three-phase non-uniform loads
To address this issue, based on a modified genetic algorithm, a novel data-driven phase identification approach is introduced to identify the phase labels under incomplete datasets [jimenez2021new]. Based on incomplete and asynchronous SM data, an optimization-based data-driven approach is constructed to recover the approximate topologies in LVDNs with open datasets [liu2025topology]. Voltage magnitude fluctuations among customers connected to different phases may exhibit similarities due to variations induced by the medium-voltage (MV) network or similar load patterns. To mitigate this effect, a high-pass filter is applied during time-series data pre-processing in the phase identification approach [hosseini2020machine]. Additionally, to mitigate the impact of electrical distance on the clustering of SM time-series data, a power-band-based strategy is proposed to select data sets, thereby improving the precision of phase identification and transformer-to-meter matching [10533846]. An optimisation model-based approach was introduced in [10742905] to estimate customer phase labels and switch states in LVDNs with limited available recordings. SA strategies were integrated to extract underlying features, thereby enhancing the accuracy of phase-label identification [7604110]. Nevertheless, the impacts of household DERs have not been sufficiently addressed in the above research. With the increasing penetration of household PV installations, the similarity among household voltage magnitudes is expected to increase significantly. Specifically, within the same LVDN, PV systems are exposed to similar solar irradiance conditions, resulting in highly correlated PV power generation patterns. This leads to similar voltage magnitude distributions among households, challenging the feasibility of correlation analysis-based approaches.
Topology identification generally aims to determine the structure of an LVDN when the network topology is completely unknown. In this setting, the potentially valuable information contained in previously recorded topologies is typically not utilised. In contrast, topology correction focuses on updating only the incorrect or outdated components of an existing topology by leveraging limited recent data while fully exploiting the information available in historical topology records. While topology identification has attracted significant research attention in recent years, topology correction remains relatively underexplored. To dynamically update the connection and the parameters after topology configuration, a graph fusion network-based approach presented in [10734663] is used to extract correct connection lines. However, the adjacency matrix is also required. For networks with low observability, a distance matrix is introduced in [10612997] to correct the topology and is also effective in detecting latent nodes positioned between observed nodes along network paths (i.e., intermediate latent nodes rather than endpoints). A physics-informed neural network is constructed to estimate the topology by representing the nodes without meters, whose parameters are adjusted by a reinforcement learning-based approach [10138375]. A data-driven approach was introduced in [7098424] to correct the connection between meters and transformers in GIS systems using voltage datasets. A similar work in [wang2023data] established a data-driven model based on the impedance matrix to correct topology and identify the impedance of connection lines using SM data. A data-driven approach in [11063399] is to infer latent nodes and line parameters for outdated topologies using power and voltage magnitude profiles. Nevertheless, different approaches often rely on different assumptions and require different types of input SM data. As a result, DSOs need to prepare multiple datasets to perform separate identification tasks.
Contributions
@@
To overcome these limitations, we propose a data-driven approach for topology correction in LVDNs using incomplete voltage data. Table 1 highlights the key differences between the proposed approach and existing approaches. By relying solely on voltage magnitude data, this approach preserves privacy and minimizes SM updates, as it releases the need for phase angle measurements from households. The main contributions are summarized as follows:
  • Table 1: Comparison of representative topology identification and correction research in LVDNs.
    Method Ref. Input Data Robustness Output
    $V$ $P/I$ Limited GIS DERs SM Missed Switch user-feeder Phase latent
    topology data impact errors data state connections labels nodes
    ANN [9758816] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\times$
    [10138375] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$
    [10734663] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
    Linear regression [8456535] $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
    [cunha2020automated] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
    Optimization based [liu2025topology] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$
    [9409966] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
    [11063399] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
    clustering-based [8955915] $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\times$
    [10533846] $\checkmark$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
    [7098424] $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$ $\times$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\times$
    [wang2023data] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$
    Heuristic Algorithms [jimenez2021new] $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$
    [10612997] $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$
    Our Method $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\times$
    Note: Topology reconfiguration studies are not included in this comparison, as they assume a known network topology and focus on optimal network operation (e.g., loss reduction and voltage regulation), rather than topology identification or correction.
@@
  • A systematic data-driven topology correction framework is proposed to update outdated LVDN topology records through three sequential steps: (1) switch state identification, (2) user–feeder connection correction, and (3) phase label identification. The proposed framework relies solely on voltage magnitude measurements from SMs, enabling the identification and correction of multiple topology components using a unified input dataset. Compared with conventional approaches that require complete SM datasets or different types of measurements for separate tasks, the proposed method allows DSOs to correct topology information using limited and noisy SM voltage data, even when household measurements are incomplete.
@@
  • A switch state identification strategy is developed by encoding the operational states of switch bars (including multiple switchgears) using label encoding rules. In the topology correction context, historical topology records and network operation logs provide labelled samples corresponding to different switch configurations. This formulation converts the identification of multiple switchgear states into a supervised classification problem that can be efficiently solved using existing machine learning algorithms. Furthermore, a modified distance function is incorporated into traditional hierarchical clustering (HC)–based correlation analysis to enhance the distinction between geographically close loads belonging to different phases or cables, thereby improving the accuracy of user–feeder and phase identification.
  • To mitigate the influence of PV power on voltage-based correlation analysis, a time-based SM data selection strategy is proposed to extract voltage measurements during periods when household loads dominate voltage variations, and PV generation remains lower than the local demand. If PV generation forecasts are available, the selection can be refined by identifying intervals with low PV injection. Otherwise, voltage data after sunset are used as a practical utilization. This strategy reduces the distortion of correlation patterns caused by PV generation and improves the accuracy of topology correction without requiring explicit PV modeling.
@@
The proposed topology correction approach is illustrated in Fig. 2, which is composed of three steps: (1) switch state identification, (2) user-feeder connection identification and (3) phase label identification. An illustrative example is presented in Fig. 5, which shows the outputs of each step. The process is triggered when the latest voltage magnitude measurements become available or an outdated topology is detected. The input includes customer voltage magnitude and incomplete recordings of the user-feeder connection, and the output is the corrected topology with accurate customer’s user feeder connections and phase labels, including single-phase and three-phase customers.
@@
Figure 3: Illustrative example of switch deployment and states in LVDNs: 1 represents the activated switch, and vice versa [abb2025].
@@
    $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\hat{y}=\arg\max_{j}\sum_{b=1}^{B}\mathbf{I}(h_{b}(x)=j),$} (5)
where $h_{b}(x)$ is the prediction of the $b$-th tree and $\mathbf{I}(\cdot)$ is an indicator function that equals 1 if the condition is met and 0 otherwise. Given a voltage sample $V_{i}$, each trained decision tree independently predicts $h_{b}(x)$ and the final labels of the switch are obtained by the majority voting mechanism.
In the topology correction context, the use of supervised learning for switch state identification is practical. Unlike topology identification problems that assume a completely unknown network topology, topology correction relies on existing topology records and historical operation logs maintained by DSOs. Although these records may contain outdated information, they still provide useful prior knowledge about typical switch configurations. Historical voltage data collected under known operating conditions can therefore be associated with these configurations to form labelled samples for model training. Moreover, the number of feasible switch configurations within a feeder is usually limited. For example, in a switch cabinet with three switches, only 7 possible configurations exist, as shown in right part of Fig. 3. Therefore, supervised learning provides a feasible solution for identifying switch states within the proposed topology correction framework.
@@
Given the switch state obtained from Section 2.1, the feeder-transformer connection is available, and the number of feeders connected to the same transformer is assumed to be accessible for the DSO and taken as the number of clusters (i.e., $K_{s}$). Lines 1-7 in Algorithm 1 are to obtain the MFP of the input voltage magnitude dataset. Lines 8-12 are to find the link matrix of matrix $D$. From line 13 onward, the input data are clustered into a $K_{s}$ cluster using the obtained linkage matrix. If two users share the same label, they are connected to the same underground feeder, and vice versa.
@@
$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\hat{y}=\arg\max_{c}\sum_{n\in\mathcal{N}_{k}}\mathbf{I}(y_{n}=c)$} (10)
@@
Although various signal processing techniques, such as SA and high-pass filters, can be employed to mitigate high-frequency or low-frequency noise induced by household PV, the determination of appropriate parameters must be tailored to specific datasets. This requirement presents practical challenges in implementing these methods effectively. Inspired by the power-band based strategies in [10533846], a time-based data selection strategy is introduced to extract voltage measurements during periods when household loads dominate voltage variations, and PV generation remains lower than the local demand. The proposed time-based data selection strategy is easily deployed and fast. Two strategies for selecting relevant data are introduced:
@@
The proposed approach adaptively selects between Eq. (13) and Eq. (14) based on PV data availability. When PV forecasts are accessible, Eq. (14) is adopted to refine the selection by identifying intervals with low PV injection. Conversely, in the absence of such data, Eq. (13) serves as an easy and alternative by utilizing post-sunset voltage measurements, ensuring the model’s applicability. By applying these time-based selection strategies, the impact of PV-induced voltage fluctuations can be mitigated, thereby improving the robustness of correlation-based phase identification methods.
For three-phase users, SMs are assumed to provide three independent voltage time-series corresponding to each phase. In the proposed methodology, each phase is modelled as an individual dataset and processed analogously to single-phase users. The algorithm performs independent analysis on the three phase-specific time-series and subsequently assigns an ordered set of phase labels, corresponding to one of the six feasible phase permutations (A–B–C, A–C–B, B–A–C, B–C–A, C–A–B, C–B–A). This utilisation way of three-phase datasets enables consistency with the single-phase data processing, avoiding inducing extra calculation to the approach.
@@
To validate the proposed framework, this section conducts comprehensive evaluations across diverse Dutch grid scenarios. Unlike existing methods that often target isolated topology components, our approach provides an integrated solution for correcting switch states, user–feeder connections, and phase labels. Consequently, each step algorithm is benchmarked against representative machine learning techniques to demonstrate its feasibility and robustness. Furthermore, the following cases also evaluate the method’s practical feasibility for DSO deployment under varying network scales and data distributions.
The code used in this paper is available online  [^1] https://github.com/distributionnetworksTUDelft/Data-driven-approach-for-topology-correction-in-LVDNs-. The resolution of the dataset is 15 minutes, resulting in 96 data points per day. The base three-phase voltage level is 0.4 kV, and the loads of three-phase customers are non-uniformly distributed across Phases A, B, and C. The time thresholds $t_{1}$ and $t_{2}$ are set to 20 and 88. To ensure data privacy, SM datasets provided by the DSO are anonymized before analysis. The original customer identifiers and location information are removed, and the SM data are randomly assigned to end-users within the studied networks. Then, the voltage magnitude profiles are generated using a power flow model, being solved by using the Power Grid Model package [PowerGridModel].
@@
Figure 7: Switch state identification accuracy under incomplete dataset by: a) RF and b) SVM.
Table 2: Accuracy (%) comparison under multiple scenarios.
@@
Table 3: Quantitative description of test LVDN topologies.
Network Feeder number Nodes per feeder Structure type
[…]
Extended cases 3 20–30 Mixed (Simple/Complex)
@@
Given datasets with random missed data points, the accuracy of RF and SVM were assessed, and the average accuracy under five simulations is shown in Fig. 7. In general, the accuracy of the two approaches improves significantly as the number of available SMs increases and slightly decreases as the missing percentage increases, demonstrating the robustness of the proposed method with sufficient data availability. When only 2 SMs data are available, the accuracy remains consistently low (around 55%–60%), indicating the method’s limited capability under sparse data conditions. Besides, the sensitivity of SVM accuracy to the available data is less than that of RF, since the accuracy of SVM remains close when the number of available SM is larger than 10. These findings highlight the robustness of SVM and RF for switch state identification tasks in LVDNs.
@@
Table 4: Accuracy (%) comparison under multiple scenarios.